# Chinese Academy of Sciences, Beijing 100190, P. R. China

### v1

@@ 5.3 Domain wall induced density perturbation
$\sigma$ $=$ $\frac{f_{w}}{\sqrt{N}}=\frac{32\pi G\mathcal{A}\sigma_{w}t}{3\sqrt{N}}=\frac{32\pi G^{2}M_{H}\mathcal{A}\sigma_{w}}{3\sqrt{N}}$ (46)
@@
$M_{H}\lesssim 5.52\times 10^{2}M_{\odot}\left(\frac{10^{5}\leavevmode\nobreak\ \rm GeV}{\sigma_{w}^{1/3}}\right)^{3}.$ (47)
@@
In Fig. 2, we plot the variance (46) and PBH relic abundance as a function of the PBH mass with the relation $M=\gamma_{p}M_{H}$ and assuming the domain wall number $N=1$. The blue and green solid curves represent the results with the wall surface tension $\sigma_{w}^{1/3}=10^{5}$ GeV and $10^{6}$ GeV, respectively. As shown in the figure, the PBH relic abundance from the domain wall number fluctuations becomes significant when the variance $\sigma(=f_{w}/\sqrt{N})\sim 10^{-1}$, which would correspond to a PBH mass of $M\sim 10^{1}\leavevmode\nobreak\ M_{\odot}$ and $10^{-2}\leavevmode\nobreak\ M_{\odot}$ for $\sigma_{w}^{1/3}=10^{5}$ GeV and $10^{6}$ GeV, respectively, as shown in the right panel of the figure. The PBH abundance sharply increases with the increasing of domain wall energy fraction $f_{w}$ and reaches its maximum at $f_{w}=1$. For $f_{w}\simeq 1$, the PBH energy fraction reaches its maximum $\beta(M)\simeq 2\times 10^{-1}$ (see the right plot of Fig. 10 in appendix D). Beyond the maximum point, $\Omega_{\rm PBH}$ scales as $M^{-1/2}$. However, we note that the results with $f_{w}\gtrsim 1$ are not reliable. This is because the domain wall energy density dominates the radiation and the universe enters into a solid-like phase with the equation of state $\omega\simeq-2/3$. In this phase, the universe would undergo an inflation stage before the annihilation of the domain walls. In this work, we assume a radiation-dominated universe with $f_{w}\lesssim 1$ and, therefore, do not give a further discussion on this scenario.
@@ 6 Effective equation of state
The critical collapse scenario states that the horizon-size perturbations could stop expanding and collapse to form PBHs when they enter into the horizon with a density contrast exceeding a critical threshold $\delta_{c}$. Therefore, the exact value of the critical density contrast $\delta_{c}$, which depends on the equation of state $\omega$, is nontrivial for the PBH formation. Effects of $\omega$ on PBH formation have been investigated in the context of first-order phase transition Jedamzik:1999am ; Jedamzik:2024wtq and QCD phase transition Byrnes:2018clq . The main force to resist the collapse of an overdense region comes from the gas pressure, which is reduced with the decrease of the equation of state during a phase transition. Therefore, the probability of PBH formation increases with the decrease of $\omega$. Eq. (6) shows that domain walls with $\overline{v}_{w}\simeq 0.4$ in the scaling regime have a negative equation of state, $\omega\simeq-2/3$. This is because domain walls have a negative surface pressure for $\overline{v}_{w}\lesssim 0.8$ (corresponding to $\overline{v}_{w}^{2}\gamma^{2}\simeq 2$), as shown by Eq. (4). We define an effective equation of state for the mixture gas of radiation and domain walls:
$\omega_{\rm eff}=\frac{p_{w}+p_{\rm rad}}{\rho_{w}+\rho_{\rm rad}}=\frac{1-2f_{w}}{3(1+f_{w})}.$ (48)
The cosmic equation of state can get smaller when the domain walls become a nontrivial component of the universe. The variation of the critical density contrast $\delta_{c}$ as a function of the equation of state has been calculated in Ref. Musco:2012au . We find these simulation results can be reproduced by the following fit function
$\delta_{c}=\frac{A_{0}}{\left[1+(d_{0}/\omega_{\rm eff})^{s_{1}}\right]^{s_{2}}},$ (49)
@@
Figure 3: The solid curves represent the fit to the simulation results, denoted by the colored points, given in Ref. Musco:2012au with Eq. (49).
Parameter $A_{0}$ $d_{0}$ $s_{1}$ $s_{2}$
@@
Table 1: Fit results of simulations in Ref. Musco:2012au with the fit formula (49) for different values of $\alpha$.
@@
In the left plot of Fig. 4, we show the critical density contrast $\delta_{c}$ in Eq. (49) as a function of the PBH mass. We observe that $\delta_{c}$ sharply decreases when $M\sim 10^{1}\leavevmode\nobreak\ M_{\odot}$ (blue curve with $\sigma_{w}^{1/3}=10^{5}$ GeV) and $M\sim 10^{-2}\leavevmode\nobreak\ M_{\odot}$ (green curve with $\sigma_{w}^{1/3}=10^{6}$ GeV), corresponding to $f_{w}\sim 10^{-1}$. This leads to an enormous improvement in the PBH formation probability, as shown in the right plot of Fig. 4, where we adopt a constant variance $\sigma^{2}=0.004$. Therefore, once a fraction $f_{w}\sim 10^{-1}$ of the cosmic energy density is composed of the domain walls, the softened equation of state by the domain wall could also lead to a significant enhancement of the PBH relic abundance.
@@ 7 Gravitational waves from domain wall annihilation
$T_{\rm ann}\simeq 334.7\leavevmode\nobreak\ {\rm MeV}\left(\frac{1}{\mathcal{A}}\right)^{1/2}\left(\frac{10.75}{g_{*}(T_{\rm ann})}\right)^{1/4}\left(\frac{V_{\rm bias}^{1/4}}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)^{2}\left(\frac{10^{5}\leavevmode\nobreak\ {\rm GeV}}{\sigma_{w}^{1/3}}\right)^{3/2}.$ (50)
@@
$M_{H}(T_{\rm ann})\simeq\frac{\mathcal{A}\sigma_{w}}{GV_{\rm bias}}=1.34\leavevmode\nobreak\ M_{\odot}\left(\frac{\mathcal{A}}{1}\right)\left(\frac{0.1\leavevmode\nobreak\ {\rm GeV}}{V_{\rm bias}^{1/4}}\right)^{4}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ {\rm GeV}}\right)^{3}.$ (51)
@@
$f_{\rm peak}(t_{0})\simeq 1.21\times 10^{-8}\mathrm{\leavevmode\nobreak\ Hz}\left(\frac{g_{*}\left(T_{\mathrm{ann}}\right)}{10.75}\right)^{1/2}\left(\frac{g_{*s}\left(T_{\mathrm{ann}}\right)}{10.75}\right)^{-1/3}\left(\frac{T_{\mathrm{ann}}}{0.1\leavevmode\nobreak\ \rm{GeV}}\right).$ (52)
@@
$\Omega_{\mathrm{GW}}\left(f_{\mathrm{peak}}\left(t_{\mathrm{ann}}\right)\right)\simeq\frac{8\pi\tilde{\epsilon}_{\mathrm{gw}}G^{2}\mathcal{A}^{2}\sigma_{w}^{2}}{3H^{2}(t_{\rm ann})},$ (53)
with $\tilde{\epsilon}_{\rm{gw}}\simeq 0.7\pm 0.4$ Hiramatsu:2013qaa . The peak gravitational wave amplitude today is estimated as
$h^{2}\Omega_{\rm{GW}}^{\rm peak}\left(t_{0}\right)=2.4\times 10^{-10}\left(\frac{\tilde{\epsilon}_{\rm{gw}}}{0.7}\right)\left(\frac{\mathcal{A}}{1}\right)^{2}\left(\frac{g_{*,s}(T_{\rm ann})}{10.75}\right)^{-4/3}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ {\rm GeV}}\right)^{6}\left(\frac{T_{\rm ann}}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)^{-4}.$ (54)
Using Eq. (45) we have
$\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ {\rm GeV}}\right)^{3}=39.6f_{w}(T)\left(\frac{1}{\mathcal{A}}\right)\left(\frac{g_{*}(T)}{10.75}\right)^{1/2}\left(\frac{T}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)^{2}.$ (55)
Combining Eqs. (54) and (55), we have
$h^{2}\Omega_{\rm{GW}}^{\rm peak}\left(t_{0}\right)=3.8\times 10^{-9}\left(\frac{\tilde{\epsilon}_{\rm{gw}}}{0.7}\right)\left(\frac{g_{*,s}(T_{\rm ann})}{10.75}\right)^{-4/3}\left(\frac{g_{*}(T_{\rm ann})}{10.75}\right)\left(\frac{f_{w}(T_{\rm ann})}{0.1}\right)^{2}.$ (56)
The causality requires the gravitational wave spectrum scale as $\propto(f/f_{\rm peak})^{3}$ for the frequency below $f_{\rm peak}$ Caprini:2009fx . On the other hand, for $f\gtrsim f_{\rm peak}$, numerical simulations Hiramatsu:2013qaa indicate a scaling behavior $\propto(f/f_{\rm peak})^{-1}$. We adopt the following fit formula for the gravitational wave spectrum from the domain wall annihilation Ferreira:2022zzo
$h^{2}\Omega_{\rm{GW}}=h^{2}\Omega_{\rm{GW}}^{\rm peak}\frac{(\gamma_{c}+c_{1})^{c_{2}}}{(c_{1}x^{-\gamma_{c}/c_{2}}+\gamma_{c}x^{c_{1}/c_{2}})^{c_{2}}},$ (57)
where $x=f/f_{\rm peak}$, $\gamma_{c}=3$ from the requirement of causality, and $c_{1},\leavevmode\nobreak\ c_{2}\simeq 1$ from the simulations. However, $c_{1}$ and $c_{2}$ may depend on the physics model, as indicated by the simulations Hiramatsu:2012sc .
@@ 8 PTA data analysis
The recent PTA observations from NANOGrav NANOGrav:2023gor ; NANOGrav:2023hvm , CPTA Xu:2023wog , EPTA EPTA:2023fyk , and PPTA Reardon:2023gzh provide $\sim 3\sigma$ evidence for the existence of stochastic gravitational wave background (SGWB) at the nano-Hz frequency band. Various scenarios have been proposed to interpret the SGWB signal (see e.g., Madge:2023dxc ; Bian:2023dnv ; Wu:2023hsa ; Ellis:2023oxs for a summary). The domain wall annihilation is a compelling interpretation for PTAs because the bias potential induced by the QCD phase transition can naturally lead to nano-Hz gravitational waves from the annihilation of domain walls Chiang:2020aui ; Lu:2023mcz . It has been noticed in Ref. Lu:2023mcz that the domain wall should annihilate before it dominates the Universe so that the gravitational wave signal could be loud enough for the PTA experiments. As shown in this work, this would also lead to significant PBH productions due to the domain wall number density fluctuations and the equation of state reduction. Our PBH production mechanism is different from those by the collapse of closed domain walls, which depends heavily on the annihilation process of the walls Ipser:1983db ; Ferrer:2018uiu .
For our purpose, we follow the standard Bayesian statistics to analyze the PTA data, including the IPTA-DR2 dataset IPTA:datalink and the recent NG15 dataset NANOGrav:2023gor . The timing-residual cross-power spectral density measured by the PTAs is given by
$S_{ab}(f)=\Gamma_{ab}\frac{h_{c}^{2}(f)}{12\pi^{2}f^{3}},$ (58)
@@
$\Omega_{\mathrm{GW}}(f)=\frac{2\pi^{2}f^{2}h_{c}^{2}(f)}{3H_{0}^{2}}.$ (59)
@@
We employ the package PTArcade Mitridate:2023oar , which wraps the code enterprise Ellis:2020zenodo and ceffyl Lamb:2023jls , to implement the GW signal model. The Bayesian inference is adopted here to derive information on the parameters by fitting to the pulsars’ timing residuals $\delta t$, which is given by the sum of white noise and red noise, including the pulsar-intrinsic red noise and common red noise produced by a GWB (see appendix for more detail). Furthermore, the package PTMCMCSampler Ellis:2017 is employed to perform the Markov chain Monte Carlo (MCMC) sampling. In the fit to each dataset, we separately generate $2\times 10^{6}$ MC sample points. The parameter posterior distributions are derived using GetDist Lewis:2019xzd . Following Refs. Antoniadis:2022pcn ; NANOGrav:2023gor to avoid the pulsar-intrinsic excess noise at high frequencies, only the first 13 and 14 frequencies of the IPTA-DR2 and NG15 datasets are included in our analysis. The priors of the parameters for signal model and noise are provided in table 3, together with the posteriors from the analysis presented in table 4 and Fig. 11 in appendix E. It has been previously found that the domain wall scenario can improve the fit to the NANOGrav data by a Bayesian factor $B\simeq 10$ NANOGrav:2023hvm . The purpose of our fit in this work is not to repeat the model comparison but to estimate the parameters regions of our concerns.
@@
In Fig. 5 and Fig. 6, we show the interpretations of the domain wall scenario for the PTA observation, noting that our fit based on Eqs (52) and (56) is $N$-independent. The red and blue regions represent the 2D posterior distributions of $T_{\rm ann}$ and $f_{w}(T_{\rm ann})$ by fit to the NG15 and IPTA-DR2 datasets, respectively. In addition, the fit to the observations indicates that a significant cosmic energy density consists of the domain walls. The relativistic quanta released by the domain walls before their annihilation could leave an observable imprint on the standard cosmology. In order not to spoil the observations of BBN and CMB, we require the following conditions on the domain wall scenario, depending on the specific model:
@@
    $\rho_{\nu}(t_{\rm rec})=2\times\frac{7}{8}\times\frac{\pi^{2}}{30}\left(\frac{4}{11}\right)^{4/3}T_{\rm rec}^{4},$ (60)
    where the recombination temperature $T_{\rm rec}\simeq 0.3$ eV. With Eqs. (22), (33), and (39) the effective number can be estimated as
    $\Delta N_{\rm eff}$ $=$ $\frac{\rho_{\rm DR}(t_{\rm ann})}{\rho_{\nu}(t_{\rm rec})}\left(\frac{g_{*,s}(T_{\rm rec})T_{\rm rec}^{3}}{g_{*,s}(T_{\rm ann})T_{\rm ann}^{3}}\right)^{4/3}$ (61)
@@
For case I, Fig. 5 and Fig. 6 show that to interpret PTAs, the annihilation of domain walls should take place during the QCD scale before the onset of BBN, i.e., the annihilation temperature $T_{\rm ann}$ should fall in the range of $\sim 0.01-1$ GeV. Therefore, domain walls do not affect the BBN processes if the domain wall decays mainly into standard model particles. For case II, $\Delta N_{\rm eff}$ observations can put useful constraints if the domain wall decays into dark radiation. The current CMB from Planck and BAO observations provide an upper bound on the number of extra neutrino species $\Delta N_{\rm eff}\leq 0.29$ at 95% confidence level (CL) Planck:2018vyg (the black dashed lines). The upcoming CMB experiments including Simons Observatory SimonsObservatory:2018koc and the CMB-S4 experiment CMB-S4:2022ght could further tighten this upper bound to 0.11 and 0.06, respectively. Indeed, numerical estimates have indicated that at most about $(10-20)\%$ of the wall energy $\rho_{w}$ would radiate as relativistic quanta during its evolution in the scaling regime. Therefore, constraints from extra neutrino species by Eq. (61) with $p_{\rm DR}=1$ are less conservative. We expect the observations from future Simons Observatory and CMB-S4 experiments can probe part of the parameter space in case II.
@@
In Fig. 7, we show the fit results for the IPTA-DR2 and NG15 gravitational spectra. We observe that the domain wall scenario can well explain the PTA observations while a large uncertainty remains at frequencies $f\gtrsim 10$ nHz. Note that the domain walls annihilate to gravitational radiation before BBN can also make a contribution to the extra neutrino species and affect the expansion rate of the universe. The bound on the relic abundance of gravitational radiation today is related to the effective number of neutrino species by Caprini:2018mtu
$h^{2}\Omega_{\mathrm{GW}}(t_{0})\lesssim 5.6\times 10^{-6}\Delta N_{\mathrm{eff}}.$ (62)
@@
Combining Eq. (50) and Eq. (55), we obtain the bias potential in terms of $f_{w}(T_{\rm ann})$ and $T_{\rm ann}$
$\frac{V_{\rm bias}^{1/4}}{0.1\leavevmode\nobreak\ {\rm GeV}}=1.37(f_{w}(T_{\rm ann}))^{1/4}\left(\frac{T_{\rm ann}}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)\left(\frac{g_{*}(T_{\rm ann})}{10.75}\right)^{1/4}.$ (63)
Together with Eq. (55) we can therefore transfer the fit results in the $f_{w}-T_{\rm ann}$ space to the $\sigma_{w}-V_{\rm bias}$ plane. We show these results in Fig. 8, which indicate that the typical bias potential $V_{\rm bias}^{1/4}\sim\Lambda_{\rm QCD}\sim 0.1$ GeV and the surface tension $\sigma_{w}^{1/3}\sim 10^{5}$ GeV and they present a strong correlation.

### v2

@@ 6.3 Density fluctuation from isocurvature perturbation induced by domain wall network
$\sigma$ $=$ $\frac{f_{w}}{\sqrt{N}}=\frac{32\pi G\mathcal{A}\sigma_{w}t}{3\sqrt{N}}=\frac{32\pi G^{2}M_{H}\mathcal{A}\sigma_{w}}{3\sqrt{N}}$ (6.15)
@@
$M_{H}\lesssim 5.52\times 10^{2}M_{\odot}\left(\frac{10^{5}\leavevmode\nobreak\ \rm GeV}{\sigma_{w}^{1/3}}\right)^{3}.$ (6.16)
@@
In Fig. 2, we plot the variance (6.15) and PBH relic abundance as a function of the PBH mass with the relation $M=\gamma_{p}M_{H}$ and assuming the domain wall number $N=1$. The blue and green solid curves represent the results with the wall surface tension $\sigma_{w}^{1/3}=10^{5}$ GeV and $10^{6}$ GeV, respectively. As shown in the figure, the PBH relic abundance from the domain wall number fluctuations becomes significant when the variance $\sigma(=f_{w}/\sqrt{N})\sim 10^{-1}$, which would correspond to a PBH mass of $M\sim 10^{1}\leavevmode\nobreak\ M_{\odot}$ and $10^{-2}\leavevmode\nobreak\ M_{\odot}$ for $\sigma_{w}^{1/3}=10^{5}$ GeV and $10^{6}$ GeV, respectively, as shown in the right panel of the figure. The PBH abundance sharply increases with the increasing of domain wall energy fraction $f_{w}$ and reaches its maximum at $f_{w}=1$. For $f_{w}\simeq 1$, the PBH energy fraction reaches its maximum $\beta(M)\simeq 2\times 10^{-1}$ (see the right plot of Fig. 11 in appendix D). Beyond the maximum point, $\Omega_{\rm PBH}$ scales as $M^{-1/2}$. However, we note that the results with $f_{w}\gtrsim 1$ are not reliable. This is because the domain wall energy density dominates the radiation and the Universe enters into a solid-like phase with the equation of state $\omega\simeq-2/3$. In this phase, the Universe would undergo an inflation stage before the annihilation of the domain walls. In this work, we assume a radiation-dominated Universe with $f_{w}\lesssim 1$ and, therefore, do not give a further discussion on this scenario.
@@ 7 Effective equation of state
The critical collapse scenario states that the horizon-size perturbations could stop expanding and collapse to form PBHs when they enter into the horizon with a density contrast exceeding a critical threshold $\delta_{c}$. Therefore, the exact value of the critical density contrast $\delta_{c}$, which depends on the equation of state $\omega$, is nontrivial for the PBH formation. Effects of $\omega$ on PBH formation have been investigated in the context of first-order phase transition [78, 79] and QCD phase transition [67]. The main force to resist the collapse of an overdense region comes from the gas pressure, which is reduced with the decrease of the equation of state during a phase transition. Therefore, the probability of PBH formation increases with the decrease of $\omega$. Eq. (2.6) shows that domain walls with $\overline{v}_{w}\simeq 0.4$ in the scaling regime have a negative equation of state, $\omega\simeq-2/3$. This is because domain walls have a negative surface pressure for $\overline{v}_{w}\lesssim 0.8$ (corresponding to $\overline{v}_{w}^{2}\gamma^{2}\simeq 2$), as shown by Eq. (2.4). We define an effective equation of state for the mixture gas of radiation and domain walls:
$\omega_{\rm eff}=\frac{p_{w}+p_{\rm rad}}{\rho_{w}+\rho_{\rm rad}}=\frac{1-2f_{w}}{3(1+f_{w})}.$ (7.1)
The cosmic equation of state can get smaller when the domain walls become a nontrivial component of the Universe. The variation of the critical density contrast $\delta_{c}$ as a function of the equation of state has been calculated in Ref. [63]. We find these simulation results can be reproduced by the following fit function
$\delta_{c}=\frac{\delta_{0}}{\left[1+(\omega_{0}/\omega_{\rm eff})^{s_{1}}\right]^{s_{2}}},$ (7.2)
@@
Figure 3: The solid curves represent the fit to the simulation results, denoted by the colored points, given in Ref. [63] with Eq. (7.2).
Parameter $\delta_{0}$ $\omega_{0}$ $s_{1}$ $s_{2}$
@@
Table 1: Fit results of simulations in Ref. [63] with the fit formula (7.2) for different values of $\alpha$.
@@
In the left plot of Fig. 4, we show the critical density contrast $\delta_{c}$ in Eq. (7.2) as a function of the PBH mass. We observe that $\delta_{c}$ sharply decreases when $M\sim 10^{1}\leavevmode\nobreak\ M_{\odot}$ (blue curve with $\sigma_{w}^{1/3}=10^{5}$ GeV) and $M\sim 10^{-2}\leavevmode\nobreak\ M_{\odot}$ (green curve with $\sigma_{w}^{1/3}=10^{6}$ GeV), corresponding to $f_{w}\sim 10^{-1}$. This leads to an enormous improvement in the PBH formation probability, as shown in the right plot of Fig. 4, where we adopt a constant variance $\sigma^{2}=0.004$. Therefore, once a fraction $f_{w}\sim 10^{-1}$ of the cosmic energy density is composed of the domain walls, the softened equation of state by the domain wall could also lead to a significant enhancement of the PBH relic abundance.
@@ 8 Gravitational waves from domain wall annihilation
$T_{\rm ann}\simeq 334.7\leavevmode\nobreak\ {\rm MeV}\left(\frac{1}{\mathcal{A}}\right)^{1/2}\left(\frac{10.75}{g_{*}(T_{\rm ann})}\right)^{1/4}\left(\frac{V_{\rm bias}^{1/4}}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)^{2}\left(\frac{10^{5}\leavevmode\nobreak\ {\rm GeV}}{\sigma_{w}^{1/3}}\right)^{3/2}.$ (8.1)
@@
$M_{H}(T_{\rm ann})\simeq\frac{\mathcal{A}\sigma_{w}}{GV_{\rm bias}}=1.34\leavevmode\nobreak\ M_{\odot}\left(\frac{\mathcal{A}}{1}\right)\left(\frac{0.1\leavevmode\nobreak\ {\rm GeV}}{V_{\rm bias}^{1/4}}\right)^{4}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ {\rm GeV}}\right)^{3}.$ (8.2)
@@
$f_{\rm peak}(t_{0})\simeq 1.21\times 10^{-8}\mathrm{\leavevmode\nobreak\ Hz}\left(\frac{g_{*}\left(T_{\mathrm{ann}}\right)}{10.75}\right)^{1/2}\left(\frac{g_{*s}\left(T_{\mathrm{ann}}\right)}{10.75}\right)^{-1/3}\left(\frac{T_{\mathrm{ann}}}{0.1\leavevmode\nobreak\ \rm{GeV}}\right).$ (8.3)
@@
$\Omega_{\mathrm{GW}}\left(f_{\mathrm{peak}}\left(t_{\mathrm{ann}}\right)\right)\simeq\frac{8\pi\tilde{\epsilon}_{\mathrm{gw}}G^{2}\mathcal{A}^{2}\sigma_{w}^{2}}{3H^{2}(t_{\rm ann})},$ (8.4)
with $\tilde{\epsilon}_{\rm{gw}}\simeq 0.7\pm 0.4$ [80]. The peak gravitational wave amplitude today is estimated as
$h^{2}\Omega_{\rm{GW}}^{\rm peak}\left(t_{0}\right)=2.4\times 10^{-10}\left(\frac{\tilde{\epsilon}_{\rm{gw}}}{0.7}\right)\left(\frac{\mathcal{A}}{1}\right)^{2}\left(\frac{g_{*,s}(T_{\rm ann})}{10.75}\right)^{-4/3}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ {\rm GeV}}\right)^{6}\left(\frac{T_{\rm ann}}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)^{-4}.$ (8.5)
Using Eq. (6.14) we have
$\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ {\rm GeV}}\right)^{3}=39.6f_{w}(T)\left(\frac{1}{\mathcal{A}}\right)\left(\frac{g_{*}(T)}{10.75}\right)^{1/2}\left(\frac{T}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)^{2}.$ (8.6)
Combining Eqs. (8.5) and (8.6), we have
$h^{2}\Omega_{\rm{GW}}^{\rm peak}\left(t_{0}\right)=3.8\times 10^{-9}\left(\frac{\tilde{\epsilon}_{\rm{gw}}}{0.7}\right)\left(\frac{g_{*,s}(T_{\rm ann})}{10.75}\right)^{-4/3}\left(\frac{g_{*}(T_{\rm ann})}{10.75}\right)\left(\frac{f_{w}(T_{\rm ann})}{0.1}\right)^{2}.$ (8.7)
The causality requires the gravitational wave spectrum scale as $\propto(f/f_{\rm peak})^{3}$ for the frequency below $f_{\rm peak}$ [81]. On the other hand, for $f\gtrsim f_{\rm peak}$, numerical simulations [80] indicate a scaling behavior $\propto(f/f_{\rm peak})^{-1}$. We adopt the following fit formula for the gravitational wave spectrum from the domain wall annihilation [82]
$h^{2}\Omega_{\rm{GW}}=h^{2}\Omega_{\rm{GW}}^{\rm peak}\frac{(\gamma_{c}+c_{1})^{c_{2}}}{(c_{1}x^{-\gamma_{c}/c_{2}}+\gamma_{c}x^{c_{1}/c_{2}})^{c_{2}}},$ (8.8)
where $x=f/f_{\rm peak}$, $\gamma_{c}=3$ from the requirement of causality, and $c_{1},\leavevmode\nobreak\ c_{2}\simeq 1$ from the simulations. However, $c_{1}$ and $c_{2}$ may depend on the physics model, as indicated by the simulations [55].
@@ 9 PTA data analysis
The recent PTA observations from NANOGrav [83, 84], CPTA [85], EPTA [86], and PPTA [87] provide $\sim 3\sigma$ evidence for the existence of stochastic gravitational wave background (SGWB) at the nano-Hz frequency band. Various scenarios have been proposed to interpret the SGWB signal (see e.g., [88, 89, 90, 91] for a summary). The domain wall annihilation is a compelling interpretation for PTAs because the bias potential induced by the QCD phase transition can naturally lead to nano-Hz gravitational waves from the annihilation of domain walls [30, 29]. It has been noticed in Ref. [29] that the domain wall should annihilate before it dominates the Universe so that the gravitational wave signal could be loud enough for the PTA experiments. As shown in this work, this would also lead to significant PBH productions due to the domain wall number density fluctuations and the equation of state reduction. Our PBH production mechanism is different from those by the collapse of closed domain walls, which depends heavily on the annihilation process of the walls [39, 41].
For our purpose, we follow the standard Bayesian statistics to analyze the PTA data, including the IPTA-DR2 dataset [92] and the recent NG15 dataset [83]. The timing-residual cross-power spectral density measured by the PTAs is given by
$S_{ab}(f)=\Gamma_{ab}\frac{h_{c}^{2}(f)}{12\pi^{2}f^{3}},$ (9.1)
@@
$\Omega_{\mathrm{GW}}(f)=\frac{2\pi^{2}f^{2}h_{c}^{2}(f)}{3H_{0}^{2}}.$ (9.2)
@@
We employ the package PTArcade [93], which wraps the code enterprise [94] and ceffyl [95], to implement the GW signal model. The Bayesian inference is adopted here to derive information on the parameters by fitting to the pulsars’ timing residuals $\delta t$, which is given by the sum of white noise and red noise, including the pulsar-intrinsic red noise and common red noise produced by a GWB (see appendix for more detail). Furthermore, the package PTMCMCSampler [96] is employed to perform the Markov chain Monte Carlo (MCMC) sampling. In the fit to each dataset, we separately generate $2\times 10^{6}$ MC sample points. The parameter posterior distributions are derived using GetDist [97]. Following Refs. [98, 83] to avoid the pulsar-intrinsic excess noise at high frequencies, only the first 13 and 14 frequencies of the IPTA-DR2 and NG15 datasets are included in our analysis. The priors of the parameters for signal model and noise are provided in table 3, together with the posteriors from the analysis presented in table 4 and Fig. 12 in appendix E. It has been previously found that the domain wall scenario can improve the fit to the NANOGrav data by a Bayesian factor $B\simeq 10$ [84]. The purpose of our fit in this work is not to repeat the model comparison but to estimate the parameters regions of our concerns.
@@
In Fig. 5 and Fig. 6, we show the interpretations of the domain wall scenario for the PTA observation, noting that our fit based on Eqs (8.3) and (8.7) is $N$-independent. The red and blue regions represent the 2D posterior distributions of $T_{\rm ann}$ and $f_{w}(T_{\rm ann})$ by fit to the NG15 and IPTA-DR2 datasets, respectively. In addition, the fit to the observations indicates that a significant cosmic energy density consists of the domain walls. The relativistic quanta released by the domain walls before their annihilation could leave an observable imprint on the standard cosmology. In order not to spoil the observations of BBN and CMB, we require the following conditions on the domain wall scenario, depending on the specific model:
@@
    $\rho_{\nu}(t_{\rm rec})=2\times\frac{7}{8}\times\frac{\pi^{2}}{30}\left(\frac{4}{11}\right)^{4/3}T_{\rm rec}^{4},$ (9.3)
    where the recombination temperature $T_{\rm rec}\simeq 0.3$ eV. With Eqs. (5.5), (6.3), and (6.9) the effective number can be estimated as
    $\Delta N_{\rm eff}$ $=$ $\frac{\rho_{\rm DR}(t_{\rm ann})}{\rho_{\nu}(t_{\rm rec})}\left(\frac{g_{*,s}(T_{\rm rec})T_{\rm rec}^{3}}{g_{*,s}(T_{\rm ann})T_{\rm ann}^{3}}\right)^{4/3}$ (9.4)
@@
For case I, Fig. 5 and Fig. 6 show that to interpret PTAs, the annihilation of domain walls should take place during the QCD scale before the onset of BBN, i.e., the annihilation temperature $T_{\rm ann}$ should fall in the range of $\sim 0.01-1$ GeV. Therefore, domain walls do not affect the BBN processes if the domain wall decays mainly into standard model particles. For case II, $\Delta N_{\rm eff}$ observations can put useful constraints if the domain wall decays into dark radiation. The current CMB from Planck and BAO observations provide an upper bound on the number of extra neutrino species $\Delta N_{\rm eff}\leq 0.29$ at 95% confidence level (CL) [68] (the black dashed lines). The upcoming CMB experiments including Simons Observatory [99] and the CMB-S4 experiment [100] could further tighten this upper bound to 0.11 and 0.06, respectively. Indeed, numerical estimates have indicated that at most about $(10-20)\%$ of the wall energy $\rho_{w}$ would radiate as relativistic quanta during its evolution in the scaling regime. Therefore, constraints from extra neutrino species by Eq. (9.4) with $p_{\rm DR}=1$ are less conservative. We expect the observations from future Simons Observatory and CMB-S4 experiments can probe part of the parameter space in case II.
@@
In Fig. 7, we show the fit results for the IPTA-DR2 and NG15 gravitational spectra. We observe that the domain wall scenario can well explain the PTA observations while a large uncertainty remains at frequencies $f\gtrsim 10$ nHz. Note that the domain walls annihilate to gravitational radiation before BBN can also make a contribution to the extra neutrino species and affect the expansion rate of the Universe. The bound on the relic abundance of gravitational radiation today is related to the effective number of neutrino species by [101]
$h^{2}\Omega_{\mathrm{GW}}(t_{0})\lesssim 5.6\times 10^{-6}\Delta N_{\mathrm{eff}}.$ (9.5)
@@
Combining Eq. (8.1) and Eq. (8.6), we obtain the bias potential in terms of $f_{w}(T_{\rm ann})$ and $T_{\rm ann}$
$\frac{V_{\rm bias}^{1/4}}{0.1\leavevmode\nobreak\ {\rm GeV}}=1.37(f_{w}(T_{\rm ann}))^{1/4}\left(\frac{T_{\rm ann}}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)\left(\frac{g_{*}(T_{\rm ann})}{10.75}\right)^{1/4}.$ (9.6)
Together with Eq. (8.6) we can therefore transfer the fit results in the $f_{w}-T_{\rm ann}$ space to the $\sigma_{w}-V_{\rm bias}$ plane. We show these results in Fig. 8, which indicate that the typical bias potential $V_{\rm bias}^{1/4}\sim\Lambda_{\rm QCD}\sim 0.1$ GeV and the surface tension $\sigma_{w}^{1/3}\sim 10^{5}$ GeV and they present a strong correlation.